You have identified a new species of flowering plant with red flowers instead of the usual blue flowers. You performed a cross and found that all of the F1 plants had purple flowers. When you self-fertilized these F1 plants, the resulting F2 was: Blue flowers: 1034 Purple flowers: 297 Red flowers: 94 What would the expected numbers be if these phenotypes were due to two alleles with incomplete dominance at a single loci? What if these phenotypes were due to two seperate genes demonstrating genetic interaction (duplicate interaction)? (2 marks)

b) Assuming that there are only two genes involved in generating these phenotypes, explain the type of interaction that could explain the observed results. Perform a Chi-square analysis to test your theory, including your hypothesis and clearly stating your conclusion.
Answer #1

Answer:

Dominant epistasis: Complete dominance at both gene pairs; however, when one gene is dominant, it hides the phenotype of the other gene. It is caused by the dominant allele of one gene, masking the action of either allele of the other gene.

Blue = A_B_ or A_bb = 12/16

Purple = aaB_ = 3/16

Red: aabb = 1/16

Expected numbers:

Total = 1034 + 297 + 94 =

Blue = 12/16 * 1425 = 1068.75

Purple = 3/16 * 1425 = 267.19

Red = 1/16*1425 = 89.06

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

Blue

1034

1068.75

-34.75

1207.56

1.1299

Purple

297

267.19

29.81

888.64

3.3259

Red

94

89.06

4.94

24.40

0.2740

1425

1425

4.7298

Chi-square value = 4.73

Degrees of freedom = Number of phenotypes – 1

Df = 3-1=2

Chi-square value of 4.73 is lesser than the critical value of 5.99. We can accept the null hypothesis.